using whole number of inches only, give the widths and lengths of six different rectangles that have an area of 90 square inches.
Found 2 solutions by Edwin McCravy, Leaf W.:
Answer by Edwin McCravy(20055)

9in x 10in, 6in x 15in, 5in x 18in, 3in x 30in, 2in x 45in, 1in x 90in

Answer by Leaf W.(135)

To get an area of 90 square inches, you need a width and length that multiply together for a product of 90. Therefore, you can just find two numbers that fulfill this requirement:
1*90 = 90, so you can have a length of 1 inch and a width of 90 inches (or vice versa).
2*45 = 90, so you can have a length of 2 inches and a width of 45 inches (or vice versa).
3*30 = 90, so you can have a length of 3 inches and a width of 30 inches (or vice versa).
5*18 = 90, so you can have a length of 5 inches and a width of 18 inches (or vice versa).
6*15 = 90, so you can have a length of 6 inches and a width of 15 inches (or vice versa).
9*10 = 90, so you can have a length of 9 inches and a width of 10 inches (or vice versa).
Good luck! =)